find an arithmetic sequence of four terms whose sum is 10, if sum of their squares is 70
Answer by richwmiller(17219)

S = ½(2a + d(n-1))n
10= 2(2a + 3d)
10= 4a + 6d
Try a=1 d=1
or
S = ½(a + an)n
10 = ½(1 + 4)4
10=2*5

1,2,3,4
1+4+9+16=30
2+3+4+5=14
not possible.